OpenStudy (anonymous):

Given the linear equation y=-2/3x-2 , find the y-coordinates of the points (-9,  ), (-3,  ), and (6,  ). Please show all of your work. Plot those points and graph the linear equation.

OpenStudy (mertsj):

y = -2/3x -2 if x = -9, replace x with -9 and find y y = -2/3(-9) -2 y = 6-2=4 if x = -3 y = -2/3(-3) - 2 y = 0 If x = 6 y = -2/3(6)-2 y = -4-2=-6